OpenStudy (djbreezy):

Which is a solution of the equation? 2x – y = 0 3 (2, –4) (1, 2) (3, –1)

OpenStudy (welshfella):

plug in x = 2 and y = -4 to test if its the second one if they fit thats the one Do same for the other choices.

OpenStudy (djbreezy):

I think its B

OpenStudy (djbreezy):

@Itrout

OpenStudy (djbreezy):

@ltrout

OpenStudy (welshfella):

2(2) - (-4) = 0???

OpenStudy (welshfella):

= 4 + 4 = 8 Its not B

OpenStudy (pawanyadav):

2x-y=0 2x=y x=y/2 So in 3rd option we get x =y/2. (1,2)

OpenStudy (welshfella):

@Pawanyadav - it would be better if you tried to get the poster involved in the solution than to do it for him.

OpenStudy (welshfella):

He/she would learn more that way.

OpenStudy (pawanyadav):

I'm expecting this... Sorry